OpenStudy (anonymous):

ISOSCELES TRAPEZOID jklm, LEG JK=5X+10, BASE KL=6X+2, LEG LM=2X+8, FIND VALUE OF X

Directrix (directrix):

Theorem: If a trapezoid is isosceles, then its legs are congruent by definition. JK = ML 5x + 10 = 2x + 8 3x = -2 x = -2/3 -->